OpenStudy (anonymous):

you roll two dice. What is the probability that at least one die shows a 4?

OpenStudy (anonymous):

1:6

OpenStudy (anonymous):

11/36?

OpenStudy (yash2651995):

its 11/36

OpenStudy (anonymous):

The probability of one dice shows 4 is \(\frac16\). Probability at least one shows 4 is \(\frac16+\frac16=\frac13 \)

OpenStudy (yash2651995):

i got it.. look say that no die should have 4... so favourable cases will be.. the one satisfying these condition.. die one can have 1,2,3,5,6 and die two can have 1,2,3,5,6 so total no of favourable cases= 5x5 = 25 total cases.. 36 P(E')= 25/36 P(E)=1- (25/36) =11/36

OpenStudy (anonymous):

yep I missed one again thanks for clarifying this for me

OpenStudy (anonymous):

my final answer is 11/36. k thanks

OpenStudy (yash2651995):

i'm sure my ans is clear